How many liters of a 10% alcohol solution must be mixed with 20 litres of a 90% solution to get a 60% solution?
Answer by rwm(914)

x+20=20+x
.1x+.9*20=.6(20+x)
x=12
use 12 liters of 10%
check
.1*12+.9*20=.6*32
true